solve each system of equations by using substitution
2x+y=4
3x+2y=1
my work so far was 3x+2(2x+4)=1 and then 3x+4x+8=1 and then you add the x's so 7x+8=1
Now what?

Answer by scott8148(6628)

2x+y=4 ___ y = 4 - 2x

substituting ___ 3x + 2(4 - 2x) = 1 ___ 8 - x = 1 ___ x = 7

substituting ___ y = 4 - 2(7) ___ y = -10